How to find d answer 2 a definite integral questin usin dimensional analysis?? 

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Dimensions does not matter in the case of the indefinite integral. But in case of the definite integral, we need to take care of dimension which we can do by comparing left hand side & right hand side with the help of dimensional analysis.
For, example we the finding the area
dA = dx.dy
\int dA = \int dx.dy
This integral does not make any sense. But
\int_{R} dA = \int_{R} dx.dy
This integral make sense where R is the two dimensional region & integral over R can be iterated into two integral.
\int_{R} dA = \int\int dx.dy
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty